# villa-X: Enhancing Latent Action Modeling in Vision-Language-Action Models

### v2

@@
Visual-Language-Action (VLA) models have emerged as a popular paradigm for learning robot manipulation policies that can follow language instructions and generalize to novel scenarios. Recent work has begun to explore the incorporation of latent actions, an abstract representation of visual change between two frames, into VLA pre-training. In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances latent action modeling for learning generalizable robot manipulation policies. Our approach improves both how latent actions are learned and how they are incorporated into VLA pre-training. Together, these contributions enable villa-X to achieve superior performance across simulated environments including SIMPLER and LIBERO, as well as on two real-world robot setups including gripper and dexterous hand manipulation. We believe the ViLLA paradigm holds significant promise, and that our villa-X provides a strong foundation for future research.
Keywords: Latent Action, Vision-Language-Action Model code: https://github.com/microsoft/villa-x site: https://aka.ms/villa-x
@@ 1 Introduction
Latent action learning has emerged as a promising approach for the pretraining of vision-language-action (VLA) models [5, 31, 52, 11, 30, 68, 35, 13], enabling learning from both robot data and human video data [67, 11, 52, 1]. At the core of these methods is a Latent Action Model (LAM), which compresses visual changes between successive frames into latent tokens, referred to as latent actions. These latent actions are expected to capture motion semantics. These tokens function as pseudo-action labels for imitation learning and enriching robot policy training with abundant, action-free data.
Yet two questions remain: how to learn latent actions and how to incorporate latent actions into VLA pre-training? In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances both key aspects of latent action modeling. For the latent action learning component, a key limitation is the absence of direct alignment between visual transitions and underlying robot states and actions. Although robot data with low-level states and actions are often available alongside action-free videos during pretraining, current methods choose to ignore these signals and focus solely on the visual part. As a result, latent actions can remain ungrounded in the robot’s physical dynamics, resulting in a weak correspondence between latent action and robot behavior. To address this issue, we incorporate a proprio Forward Dynamics Model (FDM) module into the LAM as an auxiliary decoder. This module predicts future robot proprioceptive states and actions, encouraging the learned latent actions to better reflect the agent’s behavior—not just the observed visual changes. Moreover, this design improves the interpretability of the learned latent actions and enables them to be more easily translated into executable robot actions. As a result, latent actions become a robust intermediary between vision-based representations and low-level controls. To exploit this structure, we propose to model the latent action and robot action distribution jointly through a joint diffusion process, and conditioning robot action generation on latent actions generation through attention. Compared to existing methods on latent action pre-training, our approach allows for more effective and structured information transfer from latent actions to robot actions.
We conduct comprehensive evaluations of villa-X on two simulated environments, SIMPLER and LIBERO, as well as two real-world setups, including various robot platforms with gripper and dexterous hand manipulation. Our results demonstrate that villa-X achieves superior performance, laying a robust foundation for future research in this domain.
Our contributions are listed as follows:
@@
  • We improve latent action learning by introducing an extra proprio FDM, which aligns latent tokens with underlying robot states and actions and grounds them in physical dynamics.
@@
  • Our method demonstrates superior performance on simulated environments as well as on real-world robotic tasks. The latent action expert can effectively plan into future with both visual and proprio state planning.
Figure 1: (a) High-level architecture of our villa-X method. (b) Detailed structure of the Latent Action Model (LAM) component within villa-X.
@@ 2 Related Work
Visual-Language-Action Models
Vision-Language-Action (VLA) models [5, 31, 52, 11, 30, 68, 35, 13] leverage pre-trained vision-language models (VLMs) to generate robot actions using visual and language cues. They either directly repurpose VLMs for action prediction [11, 67, 35, 30] or use action experts to map VLM outputs to robot actions [52, 5, 31]. While training on large-scale datasets like Open X-Embodiment [13] enhance the generalization ability of VLAs, cross-embodiment generalization remains challenging due to diverse robot setups and configurations. Utilizing unlabeled trajectory data with pseudo-labels such as latent actions [67, 11], language sub-goals [49], or visual sub-goals [68] supports overcoming these challenges. Our method adopts the latent action approach and improves both the modeling of latent actions and their integration into VLA pretraining.
@@
Recent exploration into latent actions began with LAPO [58] and Genie [7], which primarily focused on the video game domain. Dynamo [15] adopted a similar architecture, using inverse and forward dynamics models to shape state representations. LAOM [51] propose to use supervision to learn better latent actions in the presense of distractors on Mujoco environments.
For robotic learning, methods have started to incorporate latent actions into VLA pretraining [67, 10, 11, 52, 1, 8, 66, 38]. LAPA [67] proposes to learn from videos, and trains its latent actions and Vision-Language Model (VLM) using either human or robot video data. Concurrently, IGOR [10] learns latent actions from a mixture of human and robot videos, marking the first demonstration of successful latent action transfer between humans and robots in a unified action space for embodied AI. Moto-GPT [11] co-fine-tunes both latent and robot action labels. GR00T [52] treats latent actions as a distinct embodiment, while Go-1 [1] generates robot actions conditioned on discrete latent tokens. UniVLA [8] proposes a two-stage training pipeline to learn task-centric latent actions. More recent works like [66, 38] explore the continuous latent actions. By contrast, our approach jointly models latent and robot actions through a joint diffusion process, conditioning robot action generation on latent actions for more effective and structured information transfer. Our method offers better integration compared to LAPA [67] and GR00T [52], incorporates immediate visual context unlike Moto-GPT [11], and avoids inconsistencies caused by teacher forcing seen in Go-1 [1], achieving robust test-time reasoning.
@@ 3 Method
Figure 2: Visualization of image pairs with similar latent actions.
At a high level, our villa-X consists of two main components, as illustrated in Figure 1a:
@@
    a LAM (Latent Action Model) module that infers latent actions from a pair of observations;
@@
    a ACT (Actor) module that jointly models latent actions and robot actions given an initial visual observation and a textual task instruction.
We describe each component in detail in the following subsections. Learning in villa-X is carried out in three stages: (1) pretraining the LAM module, (2) pretraining the ACT module, and (3) finetuning ACT on target-embodiment data.
3.1 Latent Action Model
Modeling latent actions enables the use of abundant action-free video demonstration data and also holds the promise of improving generalization between robot and human motions [67, 11]. Building on these prior works, our LAM (depicted in Figure 1b) learns an inverse dynamics model (IDM) to extract latent actions $z_{t}$ from two adjacent video frames, $o_{t}$ and $o_{t+K}$. The interval $K$ is a configurable window size, chosen to balance perceptibility and generalizability: large enough to capture meaningful motion but small enough to maintain transferability.
Previous works [67, 11] train IDM by jointly optimizing a forward dynamics model (FDM) to predict future frames. Given the current frame $o_{t}$ and a latent action $z_{t}$, the model minimizes the discrepancy between the predicted frame $\hat{o}_{t+K}$ and the ground truth $o_{t+K}$. Although robot data with low-level states and actions are often available alongside action-free video data, these approaches completely discards such information, focusing solely on visual changes. Since our ultimate objective is to use latent actions to facilitate real robots policy learning, bridging vision-language prompts and low-level controls would be beneficial. To this end, we propose augmenting LAM pretraining with proprioceptive supervision. Specifically, since $z_{t}$ represents a higher-level action abstraction, we introduce an additional proprio FDM module that predicts future robot states and actions $K$ steps ahead given the current robot state $q_{t}$ and the latent action $z_{t}$. This alignment of high-level abstractions and physical robot dynamics enhances the suitability of latent action for robot policy learning.
However, because robot trajectories are often embodiment specific, conditioning solely on $q_{t}$ and $z_{t}$ may embed platform-dependent features in the latent action, hindering cross-embodiment transfer. To alleviate this issue, we provide the proprio FDM with an additional embodiment context input (e.g., dataset ID, control frequency), isolating transferable action representations from embodiment particulars.
[Uncaptioned image] Implementation
In practice, instead of predicting a single latent action between two frames (as Figure 1b depicts), LAM receives a sequence of $T_{\text{LAM}}$ observation frames as input and outputs $T_{\text{LAM}}-1$ latent actions between consecutive frames. The inverse dynamics model first encodes the frame sequence with a Spatial–Temporal Transformer [65], then applies value quantization to produce $T_{\text{LAM}}-1$ latent actions $z\in\mathbb{R}^{N_{z}\times D_{d}}$. The FDM is implemented using a Vision Transformer (ViT), while the proprio FDM is a two-layer MLP with dual output heads. The embodiment context input is embedded using learnable embeddings and concatenated with the robot state before being fed into the proprio FDM. For further implementation details and configuration specifics, please refer to the appendix.
🖹 Training and Inference
We train LAM on diverse data sources, including robot trajectories and human egocentric videos. For data lacking low-level state and action labels, we omit the proprio FDM branch and optimize only the FDM reconstruction loss. By leveraging this diverse data, our goal is to learn a generalizable intermediate representation for downstream policy learning. Once trained, only the IDM is required to extract latent actions from a pair of frames at inference, the FDM and proprioceptive FDM serve primarily as diagnostic visualization tools. We use a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc. For more details, please refer to Appendix A.1.
3.2 Actor Module
Figure 3: Structure of the ACT module.
In the same vein as common VLA models [30, 31, 5, 67, 52], our ACT module builds upon a pre-trained vision-language model (VLM) and is trained to model a generalizable robot control policy that outputs actions given a textual task instruction and an initial observation. With the pre-trained LAM, we are able to incorporate latent actions as an additional intermediate representation in our framework, rather than directly mapping vision and language inputs to low-level robot actions. The latent actions serve as a bridging language that connects task prompts to robot action planning.
Figure 3 provides an illustration of our ACT module, which consists of three components: VLM, ACT-latent, and ACT-robot. Given a textual task prompt $l$ and a visual observation $o_{t}$, the VLM first extracts visual and language features. Then, the ACT-latent component, trained with a diffusion objective, models a sequence of $n$ latent actions $z_{t},z_{t+K},\ldots,z_{t+(n-1)K}$, conditioned on the features from the VLM. Subsequently, the ACT-robot component, also implemented as a diffusion model, predicts a sequence of $m$ low-level robot actions, conditioned on the same visual and language features as well as the latent actions. Conditioning across components is achieved using uni-directional attention over intermediate features, as indicated by the dashed line in Figure 3. Additionally, ACT-robot can incorporate wrist camera input when available. Unlike prior work [5], which feeds wrist camera data into the VLM from the start, our design is more computationally efficient. Another consideration is that we view the wrist camera input as more embodiment-specific and less generalizable, which is why we incorporate it into ACT-robot rather than the VLM. The first two components, VLM and ACT-latent, are designed to handle the more generalizable aspects of the policy.
One might ask whether the length of the robot action sequence must match that of the latent action sequence, i.e., whether $m=nK$. In our design, this is not required. We allow different lengths for latent and robot action sequences, as long as they share the same starting time step $t$. For example, we may model 10 latent actions (corresponding to 20 robot actions if $K=2$) but choose to predict only 4 robot actions (effectively relying on the first 2 latent actions). This flexibility enables us to maximize the benefits of modeling latent actions.
We highlight three key elements of our design for incorporating latent actions:
  • Latent actions as a mid-level bridge. Previous works, including LAPA [67] and GR00T [52], treat latent actions similarly to robot actions and do not explicitly model the hierarchical structure between them. For example, GR00T simply treats latent actions as an additional action type and trains them alongside robot actions within a shared diffusion transformer. In contrast, our model treats latent actions as a distinct mid-level representation that bridges high-level vision and language prompts with low-level robot actions.
  • Explicit transfer from latent to robot actions. Unlike LAPA [67] and GR00T [52], where the transfer from latent actions to robot actions occurs only implicitly through pre-trained weight initialization, our model enables an explicit connection. The robot action diffusion process is directly conditioned on the latent action diffusion process, allowing more effective and structured information transfer from latent actions to robot actions.
  • Modeling latent actions sequences. While prior works (LAPA [67] and GO-1 [1] ) incorporate latent actions, they only model one-step predictions rather than planning over longer horizons at the latent action level. In contrast, our approach models a sequence of future latent actions, enabling structured planning at both the latent and robot action levels.
Implementation  We use a pre-trained PaliGemma model [3] as our VLM. Both ACT-latent and ACT-robot adopt a transformer architecture following [69], and are trained using flow matching [39, 42]. Wrist camera images are tokenized following the approach used in HPT [63]. Similar to the proprio FDM in LAM, we incorporate embodiment context embeddings in ACT-robot to account for embodiment-specific information. Inspired by Moto [11] and RDT [43], we apply attention masking during training: with 50% probability, we completely disable attention from latent actions to robot actions; otherwise, we randomly mask out attention from 50% of the latent actions. Further implementation details are provided in the appendix.
🖹 Training, Finetuning and Inference During training, we update all components of the model, including both the pre-trained and randomly initialized parts. The model is trained jointly using losses on both latent actions and robot actions, leveraging a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc. Finetuning follows the same procedure as training, with the exception that, for unseen embodiments, we initialize a new context embedding. For robot action inference, we perform denoising steps simultaneously in both ACT-Latent and ACT-Robot. For more details, please refer to Appendix A.2.

### v3

@@
Vision-Language-Action (VLA) models have emerged as a popular paradigm for learning robot manipulation policies that can follow language instructions and generalize to novel scenarios. Recent works have begun to explore the incorporation of latent actions, abstract representations of motion between two frames, into VLA pre-training. In this paper, we introduce villa-X, a novel Vision-Language-Latent-Action (ViLLA) framework that advances latent action modeling for learning generalizable robot manipulation policies. Our approach improves both how latent actions are learned and how they are incorporated into VLA pre-training. We demonstrate that villa-X can generate latent action plans in a zero-shot fashion, even for unseen embodiments and open-vocabulary symbolic understanding. This capability enables villa-X to achieve superior performance across diverse simulation tasks in SIMPLER and on two real-world robotic setups involving both gripper and dexterous hand manipulation. These results establish villa-X as a principled and scalable paradigm for learning generalizable robot manipulation policies. We believe it provides a strong foundation for future research.
Keywords: Latent Action, Vision-Language-Action Model
@@ 1 Introduction
Latent action learning has emerged as a promising approach for the pretraining of vision-language-action (VLA) models [12, 4, 31, 35, 10, 29, 70, 50], enabling learning from both robot data and human video data [68, 10, 50, 1]. At the core of these methods is a Latent Action Model (LAM), which is designed to capture the motion semantics between successive frames into compact latent tokens. These tokens, referred to as latent actions, serve as pseudo-action labels, enriching robot policy training by enabling imitation learning on abundant, action-free data.
While promising, the central challenge still lies in improving how latent actions can enhance robot policy learning. This motivates our investigation into two pivotal questions: how to better learn latent actions, and how to more effectively integrate them into VLA pre-training? In this paper, we introduce villa-X, a novel Vision-Language-Latent-Action (ViLLA) framework that advances both key aspects of latent action modeling. For the latent action learning component, existing latent action models [68, 10, 50, 1] typically compress latent actions based on visual signals, as shown in Figure 1 (a). However, while visual changes generally align with robot physical dynamics, certain motions, such as end-effector rotations or gripper movements, are subtle in pixel changes but critical for control. Vision-based models often pay less attention to these motions, a limitation also noted in recent work [9]. As a result, the learned latent actions remain physically ungrounded, hindering effective knowledge transfer. To address this, we move beyond purely visual signals by leveraging structural cues for physical grounding. Specifically, we integrate a proprioceptive Forward Dynamics Model (proprio FDM) as an auxiliary decoder within our Latent Action Model (LAM), as shown in Figure 1 (b). This module predicts future robot proprioceptive states and actions by including embodiment context as inputs to help distinguish heterogeneous data. As a result, the latent actions become better grounded by focusing on visual changes that align with the agent’s physical dynamics. This makes the latent action a more effective bridge between vision and control, ultimately improving knowledge transfer. This framework is general and can be extended to other cues like end-effector keypoint detection or human hand pose estimation, which we leave for future work. To better leverage learned latent actions, we introduce a novel integration strategy for VLA pre-training. villa-X models latent and robot actions within a joint diffusion framework composed of two components: a latent action expert (ACT-latent) and a robot action expert (ACT-robot) as shown in Figure 2. Within this framework, an attention mechanism conditions robot action generation on latent actions generation. Compared to existing methods, this framework facilitates a more effective and structured transfer of information.
We conduct comprehensive evaluations of villa-X across diverse environments. Our experiments yield two key findings. First, extensive ablation studies confirm that our proposed improvements to the latent action model and policy architecture outperform existing methods. Second, we show that by scaled pre-training, the latent action expert effectively plans for the future, and generalizes zero-shot to unseen embodiments and open-vocabulary symbolic icons. Collectively, villa-X achieves state-of-the-art performance in various tasks including numerous simulation tasks in SIMPLER and two real-world setups featuring various robotic platforms with both gripper and dexterous hand manipulators. This establishes a robust foundation for future research in the field. In summary, our main contributions are as follows:
Figure 1: (a) A standard Latent Action Model (LAM) learns a latent action $z_{t}$ primarily through visual reconstruction, predicting a future frame $\hat{o}_{t+K}$ from the current frame $o_{t}$ and latent action $z_{t}$. (b) Our proposed model enhances this by adding a proprio-FDM. This auxiliary module predicts future robot states $\hat{q}_{t+1:t+K}$ and actions $\hat{a}_{t:t+K-1}$ conditioned on an embodiment context $c_{e}$, enabling the latent actions to be better grounded in physical dynamics.
@@
  • We improve latent action learning by introducing an extra proprio FDM, which grounds latent actions in physical dynamics by aligning with underlying robot states and actions.
@@
  • Through scaled pretraining, our latent action expert develops strong zero-shot generalization capabilities. This enables the effective transfer of knowledge across diverse simulated environments and real-world robotic tasks, leading to superior performance.
@@ 2 Related Work
Vision-Language-Action Models
Vision-Language-Action (VLA) models [4, 31, 50, 10, 29, 70, 35, 12] leverage pre-trained vision-language models (VLMs) to generate robot actions using visual and language cues. They either directly repurpose VLMs for action prediction [10, 68, 35, 29] or use action experts to map VLM outputs to robot actions [50, 4, 31]. While training on large-scale datasets like Open X-Embodiment [12] enhance the generalization ability of VLAs, cross-embodiment generalization remains challenging due to diverse robot setups and configurations. Utilizing unlabeled trajectory data with pseudo-labels such as latent actions [68, 10], language sub-goals [47], or visual sub-goals [70] supports overcoming these challenges. Our method advances the latent action framework by enhancing both the modeling of latent actions and their integration into VLA pre-training.
@@
Recent exploration into latent actions began with LAPO [59] and Genie [6], which primarily focused on the video game domain. Dynamo [14] adopted a similar architecture, using inverse and forward dynamics models to shape state representations.
For robotic learning, methods have started to incorporate latent actions into VLA pretraining [68, 9, 10, 50, 1, 7, 66, 38]. LAPA [68] proposes to learn from videos, and trains its latent actions and Vision-Language Model (VLM) using either human or robot video data. Concurrently, IGOR [9] learns latent actions from a mixture of human and robot videos, marking the first successful transfer of latent actions between humans and robots. Moto-GPT [10] co-fine-tunes both latent and robot action labels. GR00T [50] treats latent actions as a distinct embodiment, while Go-1 [1] generates robot actions conditioned on discrete latent tokens. UniVLA [7] proposes a two-stage training pipeline to learn task-centric latent actions. More recent works like Yang et al. [66], Liang et al. [38] explore the continuous latent actions. [69] provides the analysis on the learned latent actions, while LAOM [49] uses supervision to learn latent actions that are robust to distractors in MuJoCo environments. By contrast, our approach jointly models latent and robot actions through a joint diffusion process, conditioning robot action generation on latent actions for more effective and structured information transfer. Our method improves upon prior work in several key aspects: it offers tighter integration than LAPA [68] and GR00T [50]; it incorporates immediate visual context, unlike Moto-GPT [10]; and it avoids teacher-forcing inconsistencies seen in Go-1 [1]. These advantages collectively contribute to more robust reasoning at test time.
@@ 3 Method
Figure 2: Architecture of ACT: A hierarchical policy that predicts latent action plans and conditions robot action generation on them, incorporating embodiment context and attention masking.
Our method, villa-X, learns a physically grounded latent action space and uses it to train a VLA policy. The framework has two components:
@@
    Latent Action Model (LAM) infers latent actions from a pair of observations, aligning these latent actions with robot dynamics via additional proprioceptive supervision.
@@
    ACTor Module (ACT) builds on a pre-trained vision-language backbone and jointly models sequences of latent and robot actions for improved planning and control.
Training proceeds in three stages: (i) LAM pretraining on diverse datasets, (ii) ACT pretraining with joint latent–robot modeling, and (iii) embodiment-specific finetuning.
3.1 Latent Action Model (LAM)
Latent actions provide a compact intermediate representation, enabling the use of abundant human videos and improving cross-embodiment generalization [68, 10]. Prior works typically learn a quantized codebook of latent actions via two modules: an Inverse Dynamics Model (IDM) and a visual Forward Dynamics Model (FDM). The IDM predicts a latent token $z_{t}$ from a frame pair $(o_{t},o_{t+K})$, while the FDM reconstructs the future observation $\hat{o}_{t+K}$ from $(o_{t},z_{t})$:
$z_{t}=\text{IDM}(o_{t},o_{t+K}),\quad\hat{o}_{t+K}=\text{FDM}(o_{t},z_{t}).$ (1)
This objective ensures consistency in visual change but ignores physical dynamics, producing latents that are insufficiently grounded when robot states are available.
Proprioceptive Grounding.
To address this, we introduce an additional proprioceptive Forward Dynamics Model (proprio-FDM) that predicts both future robot states and actions $K$ steps ahead, given the current state $q_{t}$ and latent $z_{t}$:
$(\hat{q}_{t+1},...,\hat{q}_{t+K},\hat{a}_{t+1},...,\hat{a}_{t+K})=\text{proprio-FDM}(q_{t},z_{t},c_{e}),$ (2)
where $c_{e}$ denotes an embodiment context described below. Optimizing visual and proprioceptive forecasting jointly encourages latent tokens to emphasize physical dynamics alongside visual changes.
Disambiguating Heterogeneous Embodiments.
Large-scale datasets mix embodiments with different morphologies and control frequencies. Naively conditioning the proprio-FDM on $(q_{t},z_{t})$ risks encoding embodiment-specific features into latents. We introduce a context vector $c_{e}$ comprising:
$c_{e}=f(\text{dataset ID, control frequency}),$ (3)
where dataset IDs are mapped to learnable embeddings, and frequencies are encoded using sinusoidal features passed through an MLP. These embeddings, concatenated with $q_{t}$, allow proprio-FDM to separate embodiment-specific dynamics while preserving latent action consistency across datasets.
The full LAM thus optimizes image reconstruction loss, proprioceptive prediction loss, and vector-quantization commitments jointly. For human video lacking proprio labels, the proprio term is omitted. Finally, we adopt the continuous vector from the VQ codebook center as our latent actions. We refer readers to Appendix A for further implementation details.
In summary, our LAM extends prior latent action models beyond compressing only visual changes to jointly modeling physical state transitions. While this work leverages robot proprioception for grounding, the framework is generic: alternative structural cues like end-effector keypoint detection or human hand pose estimation could replace low-level states, which we leave for future exploration.
3.2 Actor Module (ACT)
Our ACT module extends traditional vision-language-action (VLA) approaches by explicitly modeling both latent actions ($z_{t:t+(n-1)K}^{K}=(z_{t},z_{t+K}...,z_{t+(n-1)K}$)) and robot actions ($a_{t:t+m-1}=(a_{t},a_{t+1},...,a_{t+m-1})$) with a joint diffusion process. We factorize the policy into two conditional distributions:
$\pi(a_{t:t+m-1},z_{t:t+(n-1)K}^{K}\mid o_{t},l,q_{t},c_{e})$ $=\underbrace{\pi_{\text{robot}}\bigl(a_{t:t+m-1}\mid z_{t:t+(n-1)K}^{K},\,o_{t},l,q_{t},c_{e}\bigr)}_{\text{ACT-robot}}$ (4)
$\quad\cdot\;\underbrace{\pi_{\text{latent}}\bigl(z_{t:t+(n-1)K}^{K}\mid o_{t},\,l\bigr)}_{\text{ACT-latent}}.$
where $o_{t}$ is the observation, $l$ the language instruction, $q_{t}$ the proprioceptive state, and $c_{e}$ the embodiment context. Additionally, the low-level policy can optionally incorporate wrist camera input if available. This explicit modeling and factorization improves upon prior methods, such as LAPA [68], which rely on latent actions only through pre-trained weight initialization. By contrast, our method treats latent actions as a distinct mid-level representation that bridges high-level vision and language prompts with low-level robot actions, and allow for allowing more effective and structured information transfer from latent actions to robot actions.
Architecture.
ACT (Figure  2) comprises three experts with a blockwise causal attention mask:
  • VLM: Encodes the visual-language inputs into high-level features.
  • ACT-latent: Latent action expert that predicts latent action tokens for mid-level planning, conditioning on VLM features.
  • ACT-robot: Robot action expert that produces the low-level action chunk, conditioning on VLM features, predicted latents and additional control-specific inputs including proprioceptive states and embodiment context.
Attention Masking Strategies.
A key aspect of ACT is how we maintain a robust dependence on the latent tokens without letting the policy learn trivial shortcuts. Inspired by Moto [10] and RDT [41], we adopted the stochastic Masking strategy. During training, we stochastically mask the attention from robot actions to latent actions. In 50% of the cases, all robot-to-latent attention are masked; otherwise, 50% of the latent tokens are randomly masked. This can prevent the robot-action branch from overly relying on latent actions and leading to improved robustness. We found this design crucial in practice.
Joint Diffusion Modeling.
For implementation, villa-X models the joint distribution of the future robot actions $a_{t:t+m-1}$ and latent actions $z_{t:t+(n-1)K}^{K}$ using a conditional flow matching framework. For notational simplicity, we group these actions into a single variable $x_{t}$ and denote the conditioning inputs $(o_{t},l,q_{t},c_{e})$ as $O_{t}$. The objective is to train a network $v_{\tau}^{\theta}$ that minimizes the flow matching loss:
$L_{\tau}(\theta)=\mathbb{E}_{p(x_{t}\mid O_{t}),\,q(x_{t}^{\tau}\mid x_{t})}\left\|v_{\tau}^{\theta}(x_{t}^{\tau},O_{t})-u(x_{t}^{\tau}\mid x_{t})\right\|^{2}$ (5)
where $\tau\in[0,1]$ denotes flow matching timestep. In practice, we first sample random noise $\epsilon\sim N(0,I)$ to create a noisy target $x^{\tau}_{t}=\tau x_{t}+(1-\tau)\epsilon$. The network $v_{\tau}^{\theta}(x_{t}^{\tau},O_{t})$ is then trained to predicted the denoising vector field $u(x_{t}^{\tau}\mid x_{t})=\epsilon-x_{t}$. During training, we sample $\tau$ from beta distributions. Notably, the explicit factorization in Eq. 4 is achieved by block-wise causal attention.